OpenStudy (anonymous):

soybean meal is 16% protein, cornmeal is 8% protein hpw many pounds of each sould be mixed together in order to get 320 lb mixture that is 10% protein how many pounds of cornmeal should be in the mixture how many pounds of soybean should be in the mixture

OpenStudy (anonymous):

Well let x = quantity of soybean in lb. y = quantity of cornmeal 16x + 8y = 320 x 10% = 32 x + y = 320 Solve the two equations.

OpenStudy (anonymous):

Woops i meant 0.16 and 0.08 for the coefficients ^^

OpenStudy (anonymous):

let x for 16% y for 8% 0.16X+0.08Y= 320*0.10 X+Y= 320 X=320-Y Subtitution X=320-Y 0.16(320-Y)+0.08Y= 320*0.10 51.2 -0.16Y +0.08Y= 32 -0.16Y +0.08Y= 32-51.2 -0.08 Y = -19.2 Y =-19.2\-0.08 Y =240 X+240= 320 x= 320-240 x=80

OpenStudy (anonymous):

16% need 80 lb 8% need 240 lb

OpenStudy (anonymous):

Solve these equations 0.16x + 0.08y = 32 x + y = 320 You can solve this in many ways. EIther guassian elimination, elimination or substitution. By elimination, Multiply first eq by 100 to get 16x + 8y = 3200 -> divide by 8 -> 2x + y = 400

OpenStudy (anonymous):

Now you have 2x + y = 400 x + y = 320 Subtract first eq by the second, You get x = 400-320 = 80 You got x, now solve for y. You know that x +y = 320 so y = 320-x = 320-80 = 240

OpenStudy (anonymous):

Nancy's answer is correct.

OpenStudy (anonymous):

you guys rock im copying these down for examples thank you so much for explaining the answers

OpenStudy (anonymous):

Your welcome!